OpenStudy (anonymous):

How many different batting orders are possible if 9 people are chosen from a group of 12 on a baseball team?

OpenStudy (kirbykirby):

\[ 12P9=\frac{12!}{(12-9)!}\]

OpenStudy (kirbykirby):

It's a permutation

OpenStudy (anonymous):

Could you run me through the equation? THis is a pre test and i haven't seen this before.

OpenStudy (kirbykirby):

\(12!=12\cdot 11\cdot 10 \cdot 9 \cdot 8 \cdot 7 \cdot 6 \cdot 5 \cdot 4 \cdot 3 \cdot 2 \cdot 1\)

OpenStudy (anonymous):

So i multiply all that, And i get 479001600 then i set it equal to \[\left(\begin{matrix}479001600 \\(12-9)\end{matrix}\right)\] ?

OpenStudy (kirbykirby):

in the denominator (12-9)! = (3)! = 3!

OpenStudy (anonymous):

Then i get a number that is 159667200

OpenStudy (kirbykirby):

hm, did you do 3! = 3 * 2 * 1

OpenStudy (kirbykirby):

the \(\large !\) is very important

OpenStudy (kirbykirby):

4! = 4 * 3 * 2 * 1 7! = 7*6*5*4*3*2*1 3! = 3*2*1 etc...

OpenStudy (anonymous):

Oh right, Ok so i get 79833600 and i believe that is the answer?

OpenStudy (kirbykirby):

yes =]

OpenStudy (anonymous):

Thank you so much, That really cleared up a lot

OpenStudy (kirbykirby):

your welcome